OpenStudy (anonymous):

estimate the difference first. then subtract. simplify, if necessary. 9- 4 3/7

OpenStudy (anonymous):

what wre the two seperate problems? is this only one?

OpenStudy (dumbcow):

63/7 - 31/7 = 32/7 = 4 4/7

OpenStudy (anonymous):

it didn't post the other one. here it is 3 3/5 - 7/10

OpenStudy (dumbcow):

36/10 - 7/10 = 29/10 = 2 9/10

OpenStudy (anonymous):

Your answer is 4 and 3/7

OpenStudy (anonymous):

um yeah......?????? (crickets)

OpenStudy (anonymous):

What you do is take 9 and convert it

OpenStudy (anonymous):

But first, take the 4 and 3/7ths and make it a mixed number

OpenStudy (anonymous):

This is how you do that...

OpenStudy (anonymous):

ok. im working out the problems also, n i see how both of you came up with the answer.

OpenStudy (anonymous):

You multiply the 7 in 3/7 time the 4 first

OpenStudy (anonymous):

Okay, I hope you have everything go well :D

OpenStudy (anonymous):

okay thanks if i need anymore help. i will let you know thanks.. im trying to figure it now.

OpenStudy (anonymous):

still (crickets)

OpenStudy (anonymous):

can i have a metal?

OpenStudy (anonymous):

i thought it gave it to you.. ma b got u

OpenStudy (anonymous):

thank you :) <3

OpenStudy (anonymous):

i have a ???